ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 89, 1995

Burton, R. and J. E. Steif New results on measures of maximal entropy

275

Fonf, V. P. Boundedly complete basic sequences, c_{0}-subspaces, and injections of Banach spaces

173

Gonzalo, R. and J. A. Jaramillo Smoothness and estimates of sequences in Banach spaces

321

Grabarnik, G. and E. Hrushovski Singular compactness and the Neveu decomposition

135

Haxell, P. E. and Y. Kohayakawa The size-Ramsey number of trees

261

Hrushovski, E. See G. Grabarnik

Jacquet, H. The continuous spectrum of the relative trace formula for GL(3) over a quadratic extension

1

Jara, P. and E. Santos Completions of commutative rings and modules

141

Jaramillo, J. A. See R. Gonzalo

Jiang, Q. and L. Peng Phase space, wavelet transform and Toeplitz– Hankel type operators

157

Kaimanovich, V. The Poisson boundary of covering Markov operators

77

Kechris, A. S. and S. Solecki Approximation of analytic by Borel sets and definable countable chain conditions

343

Kifer, Y. Spectrum, harmonic functions, and hyperbolic metric spaces

377

Kohayakawa, Y. See P. E. Haxell

Komorowski, R. A. and N. Tomczak-Jaegermann Banach spaces without local unconditional structure

205

Mekler, A. H. and S. Shelah Almost free algebras  237

Peng, L. See Q. Jiang

Rudelson, M. Estimates of the weak distance between finitedimensional Banach spaces

189

Santos, E. See P. Jara

Shelah, S. Every null-additive set is meager-additive  357

Shelah, S. See A. H. Mekler Solecki, S. See A. S. Kechris Srebro, U. Carathéodory balls and norm balls in H_{n} =

{z ∈ ℂ^{n}: ∥z∥_{1} < 1} 61

Steif, J. E. See R. Burton Stein, Y. The Jacobian problem as a system of ordinary differential

equations

301

Tomczak-Jaegermann, N. See R. A. Komorowski Venkataramana, T. N. On some rigid subgroups of semisimple Lie

groups

227

Zwonek, W. Carathéodory balls and norm balls of the domains H_{n} =

{z ∈ ℂ^{n}: |z_{1}| + ⋯ + |z_{n}| < 1} 71